find voltage across the resistors

+ Amp { RG=lar R2=60 m + 121 R3} SR₂-240 - 30 what is the Vollage across Resistor R3 30
Answer #1

-12 6 si M 24 30 De Current through 1201 is 1A. ... voltage ( VIR) across 12x is = 12X1 =12V There isa Voltage source acrossvoltage across 61 = 6 (11-12) 6C 6 +1) 6 (T6) 7V voltage across 301 = 30C1) =30(1/6) = 5V , . Vor=70 V 301 - SV V24л = 12V Vi